A school principal claims that grade 11 students have a mean grade of 86 with a standard deviation of 4. suppose that the distribution is approximately normal. What is the probability that a random selected grade will be less than 84?


1
Expert's answer
2022-06-01T11:46:49-0400
"P(X<84)=P(Z<\dfrac{84-86}{4})"

"=P(Z<-0.5)\approx0.3085"